• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Feuerbach

    Log entry time 09:03:09 on August 19, 2006

    Entry number 180884

    This entry is a followup to: 180635

    keyword=Summary of FPP activity late yesterday

    In the morning it was determined the gas for the FPP was "bad", we thought due to a nearly-empty dewar of Argon. After 4hrs of flushing with the new Argon bottle the efficiency had not improved, and the VDC's (which pull from the same Argon source) continued to look fine, leading us to suspect the gas mixture. Scott and Jack (remotely) investigated the source of the problem, and determined the gas mixing controller for the FPP might have a problem, and it was repaired/swapped out around 4pm.

    To monitor the chambers recovery, the HV on the chambers was increased until their efficiency was "normal". Production runs were taken during the swing and owl shift, and the HV slowly dropped over the night as the gas recovered. This morning the high voltage on the chambers are 100 to 125V higher than their nominal 1775V, and they are operating well.